I'm having problems doing this can you help me.
Volleyball teams receive 2 points when they win and 1 point when they tie. One season a volleyball team won a championship with 57 points. the team won 12 more games than they tied. How many wins did the team have? How many ties did the team have?

Answer by Alan3354(69443)

Volleyball teams receive 2 points when they win and 1 point when they tie. One season a volleyball team won a championship with 57 points. the team won 12 more games than they tied. How many wins did the team have? How many ties did the team have?
----------------
t = ties
w = wins
w = t+12 (won 12 more than tied)
t + 2w = 57
-----------
Sub for w
t + 2(t+12) = 57
t + 2t+24 = 57
3t = 33
t = 11
w = 23